OpenStudy (babyslapmafro):

How do you perform tabular integration if both functions infinitely repeat differentiation and integration respectively? (click to see example problem). Thanks

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\sin(x)\cos(x)dx\]

OpenStudy (anonymous):

Use the trigonometric identity: 2*sin(x)*cos(x) = sin(2x) http://www.sosmath.com/trig/Trig5/trig5/trig5.html Then differentiate that way. But if you need to integrate forcibly see this solution: http://answers.yahoo.com/question/index?qid=20071226235101AAaNJsH